# Simple homotopy types of even dimensional manifolds

### v2

@@ 11. Proof of main results on the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$
Let $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$. If $m\neq 15$, then Lemma 11.8 implies that $D(\mathbb{Z}C_{m})=0$ and so |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$. If $m=15$, then it is shown in [CN72, p. 48] that $|D(\mathbb{Z}C_{15})|=2$. This implies that $D(\mathbb{Z}C_{15})\cong\mathbb{Z}/2$ has the trivial action and so |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{15})\}|=1$.
By Proposition 11.6, the remaining $m\geq 2$ square-free for which |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$ are:
@@
By Proposition 10.17, we have that $\prod_{d\mid m}\operatorname{odd}(|\widetilde{V}_{d}|)$ divides |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|$, where $\widetilde{V}_{d}$ is as defined in Definition 10.15. It therefore suffices to prove that, for each $m$ listed above, we have $\operatorname{odd}(|\widetilde{V}_{d}|)\neq 1$ for some $d\mid m$.
@@
In each case, $\operatorname{odd}(c_{2p})\neq 1$ and so $\operatorname{odd}(|\widetilde{V}_{2p}|)\neq 1$. Hence |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|\neq 1$ for $m=2p$ where $p\in\{11,13,17,19,29\}$.
@@
In the cases $(p,q)\in\{(3,11),(5,7),(3,13)\}$, $\operatorname{odd}(c_{pq})\neq 1$ and so $\operatorname{odd}(|\widetilde{V}_{pq}|)\neq 1$. Hence |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|\neq 1$ for $m=pq$ or $2pq$ where $(p,q)\in\{(3,11),(5,7),(3,13)\}$.
We will deal the the three remaining cases $m=21,30$ and $42$ directly from the definition of $\widetilde{V}_{m}$:
@@
and $E=\langle\zeta_{15},\mathbb{Z}[\lambda_{15}]^{\times}\rangle$ has index two in $\mathbb{Z}[\zeta_{15}]^{\times}$. By the same argument as Proposition 10.22, we get that $|\widetilde{V}_{30}|=|\operatorname{coker}(\Psi_{30}^{+})|$ is divisible by
@@
since $f_{2}=\operatorname{ord}_{15}(2)=4$ and $g_{2}=1$. Since $\operatorname{odd}(c_{30})\neq 1$, this implies that $\operatorname{odd}(|\widetilde{V}_{30}|)\neq 1$. Hence |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{15})\}|\neq 1$.
We will deal with the remaining cases $m=21$, $42$ by computing the involution on $\widetilde{V}_{21}$ explicitly. This turns out to be necessary since, by Remark 11.12, we have $|D(\mathbb{Z}C_{m})|=4$ in each case and so $\operatorname{odd}(|\widetilde{V}_{21}|)=\operatorname{odd}(|\widetilde{V}_{42}|)=1$. If |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$ for $m=21$ or $42$ then, by Proposition 10.17, we would have $|2\cdot\widetilde{V}_{21}|=1$. Hence it suffices to prove that $|2\cdot\widetilde{V}_{21}|\neq 1$.
We now claim that $\widetilde{V}_{21}=\operatorname{coker}(\Psi_{21}^{+})\cong\mathbb{Z}/4$, which implies that $22\cdot\widetilde{V}_{21}\cong\mathbb{Z}/$ 2. First note that, by the proof of Proposition 10.22, we have that
@@
where $211\in\mathbb{Z}/$ 21 maps to $\Psi_{21}^{+}(\zeta_{21})$ and $21\in\mathbb{Z}/$ 2 maps to $\Psi_{21}^{+}(1-\zeta_{21})$.
Since $g_{3}=1$, $\mathbb{F}_{3}[\zeta_{7}]\cong\mathbb{F}_{3}(\zeta_{7})$ is a field and $\mathbb{F}_{3}[\lambda_{7}]=\mathbb{F}_{3}[\lambda_{7}]$ is a subfield. This implies that $\frac{\mathbb{F}_{3}[\zeta_{7}]^{\times}}{\mathbb{F}_{3}[\lambda_{7}]^{\times}}\cong\frac{\mathbb{Z}/(3^{6}-1)}{\mathbb{Z}/(3^{3}-1)}\cong\mathbb{Z}/28$. We have $\mathbb{F}_{7}[\zeta_{3}]=\mathbb{Z}[x]/\langle 7,1+x+x^{3}\rangle=\mathbb{Z}[x]/\langle 7,(x-2)(x-4)\rangle\cong\mathbb{F}_{7}\times\mathbb{F}_{7}$ where $\zeta_{3}\mapsto(2,4)$ and so $\mathbb{F}_{7}[\zeta_{3}]^{\times}\cong(\mathbb{Z}/6)^{2}$. Since $\mathbb{F}_{7}[\lambda_{3}]^{\times}=\mathbb{F}_{7}^{\times}\cong\mathbb{Z}/6$, this implies that $\frac{\mathbb{F}_{7}[\zeta_{3}]^{\times}}{\mathbb{F}_{7}[\lambda_{3}]^{\times}}\cong\mathbb{Z}/6$. Hence $D\coloneqq\frac{\mathbb{F}_{3}[\zeta_{7}]^{\times}}{\mathbb{F}_{3}[\lambda_{7}]^{\times}}\oplus\frac{\mathbb{F}_{7}[\zeta_{3}]^{\times}}{\mathbb{F}_{7}[\lambda_{3}]^{\times}}\cong\mathbb{Z}/28\oplus\mathbb{Z}/6$.
Now $\Psi_{21}^{+}(\zeta_{21})=[(\zeta_{7},\zeta_{3})]$ where $[\zeta_{7}]\in\mathbb{Z}/28$ has order $7$ and $[\zeta_{3}]\in\mathbb{Z}/6$ has order $3$. This implies that $\operatorname{coker}(\mathbb{Z}/21\to D)\cong\mathbb{Z}/4\oplus\mathbb{Z}/2$. Note that $\Psi_{21}^{+}(1-\zeta_{21})=[(1-\zeta_{7},1-\zeta_{3})]$. The isomorphism $\mathbb{F}_{7}[\zeta_{3}]^{\times}/\mathbb{F}_{7}[\lambda_{3}]^{\times}\to\mathbb{Z}/6$ sends $11-\zeta_{3}\mapsto 1-2=-$ 1 and so the image of $\Psi_{21}^{+}(1-\zeta_{21})$ in $\operatorname{coker}(\mathbb{Z}/21\to D)\cong\mathbb{Z}/4\oplus\mathbb{Z}/2$ has the form $(*,-1)$. Since it has order two, this implies that $\operatorname{coker}(\Psi_{21}^{+})\cong\operatorname{coker}(\mathbb{Z}/21\oplus\mathbb{Z}/2\to D)\cong\mathbb{Z}/4$ as required. ∎
We will now prove (A3). Our approach will be to use Lemma 11.4 (iv), and will be analogous to our proof of (A2). In particular, we will begin by classifying the $m\geq 2$ square-free for which $\prod_{d\mid m}|\{x\in C(\mathbb{Z}[\zeta_{d}])\mid x=-\overline{x}\}|=1$. We will then determine the subset of these values for which |$|\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}|=1$.
@@
If $h_{m}^{-}=1$, then $h_{d}^{-}=1$ for all $d\mid m$ and so $\prod_{d\mid m}|\{x\in C(\mathbb{Z}[\zeta_{d}])\mid x=-\overline{x}\}|=1$. Conversely, if $h_{m}^{-}\neq 1$ and $m\not\in\{29,39,58,78\}$, then Proposition 11.6 implies that |$|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|\neq 1$ and so |$|\{x\in C(\mathbb{Z}[\zeta_{m}])\mid x=-\overline{x}\}|\neq 1$. It now suffices to show that, if $m\in\{29,39,58,78\}$, then |$|\{x\in C(\mathbb{Z}[\zeta_{m}])\mid x=-\overline{x}\}|\neq 1$.
@@
We will now prove the following. By Lemma 11.4 (iv), this completes the proof of (A3).
@@
If |$|\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}|=1$, then |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$. By Propositions 10.5, 11.7, LABEL:, and 11.9, the $m\geq 2$ square-free for which $\prod_{d\mid m}|\{x\in C(\mathbb{Z}[\zeta_{d}])\mid x=-\overline{x}\}|=1$ and |$|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$ are as follows:
@@
If $m=p$ for $p\leq 17$ prime or $m=2p$ for $p\in\{3,5,7\}$, then Lemma 11.8 implies that $D(\mathbb{Z}C_{m})=0$ and so |$|\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}|=1$. If $m=15$, then it is shown in [CN72, p. 48] that $|D(\mathbb{Z}C_{15})|=2$. This implies that $D(\mathbb{Z}C_{15})\cong\mathbb{Z}/2$ has the trivial action and so $\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}\cong\mathbb{Z}/2$. ∎
Finally, we will prove (A4). Note that these results are implied by computations used the proofs of Propositions 11.7 and 11.10, but we will repeat them here for the convenience of the reader.
@@
This completes the proofs of (A1)-(A4) and so, by the discussion at the start of this section, completes the proofs of Theorems 11.1 and 11.2.
@@
$A_{m}\coloneqq\frac{\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid x=-\overline{x}\}}{\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}}\cong\widehat{H}^{1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{m}))$
where the isomorphism comes from Proposition 9.5. Similarly to the proofs of Theorems 11.1 and 11.2, we will begin by noting that it now suffices to prove the following two statements.
@@
To see this, note that (B1) directly implies Theorem 11.3 (i). Next, Theorem 11.3 (ii) follows from (B2) since it implies that
$\sup_{n\leq m}|A_{n}|$ $\geq\max\{|A_{2^{\lfloor\log_{2}(m)\rfloor}}|,|A_{2^{\lfloor\log_{2}(m)\rfloor-1}}|\}\geq\sqrt{|A_{2^{\lfloor\log_{2}(m)\rfloor}}|\cdot|A_{2^{\lfloor\log_{2}(m)\rfloor-1}}|}$
$\geq 2^{2^{\lfloor\log_{2}(m)\rfloor-3}-1}\geq 2^{\frac{m}{16}-1}$
@@
We will again make use of the following short exact sequence of $\mathbb{Z}C_{2}$-modules established in Section 10.1:
@@
where $D(\mathbb{Z}C_{m})$ has the induced involution, and each $C(\mathbb{Z}[\zeta_{d}])$ has the involution induced by conjugation. Each of statements (B1) and (B2) will be proven via the following lemma.
@@
[…]
@@
    If $h_{m}$ is odd, then $A_{m}\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{m})_{(2)})$
@@
    If $|D(\mathbb{Z}C_{m})|$ is odd, then $A_{m}\cong\bigoplus_{d\mid m}\widehat{H}^{1}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{(2)})$.
@@
By Proposition 9.9, we have that $A_{m}\cong\widehat{H}^{1}(C_{2};C(\mathbb{Z}C_{m})_{(2)})$. The short exact sequence stated above induces a short exact sequence on their $2$-primary submodules:
$0\to D(\mathbb{Z}C_{m})_{(2)}\to C(\mathbb{Z}C_{m})_{(2)}\to\bigoplus_{d\mid m}C(\mathbb{Z}[\zeta_{d}])_{(2)}\to 0$
@@
$\widehat{H}^{n}(C_{2};\bigoplus_{d\mid m}C(\mathbb{Z}[\zeta_{d}])_{(2)})\cong\bigoplus_{d\mid m}\widehat{H}^{n}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{(2)})$
@@
To prove (i), suppose $h_{m}$ is odd. Then $h_{d}$ is odd for all $d\mid m$ since $h_{d}\mid h_{m}$ [Wa97, p. 205]. This implies that $C(\mathbb{Z}[\zeta_{d}])_{(2)}=0$ for all $d\mid m$ and so $\bigoplus_{d\mid m}\widehat{H}^{n}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{(2)})=0$ for all $n\in\mathbb{Z}$. The 6-periodic exact sequence then gives that $A_{m}\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{m})_{(2)})$.
To prove (ii), suppose $|D(\mathbb{Z}C_{m})|$ is odd. Then $D(\mathbb{Z}C_{m})_{(2)}=0$, $\widehat{H}^{n}(C_{2};D(\mathbb{Z}C_{m})_{(2)})=0$ for all $n$, so the 6-periodic exact sequence gives that $A_{m}\cong\bigoplus_{d\mid m}\widehat{H}^{1}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{(2)})$. If $h_{m}$ and $|D(\mathbb{Z}C_{m})|$ are odd, then $D(\mathbb{Z}C_{m})_{(2)}=0$ and so $A_{m}\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{m})_{(2)})=0$, which proves (iii). ∎
We will now prove the following which implies (B1) since $h_{3}=1$ is odd (see, for example, Proposition 10.5). Note that it also implies that $|A_{p^{n}}|=1$ for $n\geq 1$ and other primes $p\neq 3$.
@@
We will now prove (B2). By the discussion above, this completes the proof of Theorem 11.3.
@@
$A_{2^{n}}\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{2^{n}})_{(2)}).$
By Propositions 9.7 and 10.19, we have a $6$-periodic exact sequence of finite abelian groups:
[…]

### v3

@@ 11. Proof of main results on the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$
Let $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$. If $m\neq 15$, then Lemma 11.7 implies that $D(\mathbb{Z}C_{m})=0$ and so $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$. If $m=15$, then it is shown in [CN72, p. 48] that $|D(\mathbb{Z}C_{15})|=2$. This implies that $D(\mathbb{Z}C_{15})\cong\mathbb{Z}/2$ has the trivial action and so $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{15})\}|=1$.
By Proposition 11.5, the remaining $m\geq 2$ square-free for which $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|=1$ are:
@@
By Proposition 10.18, we have that $\prod_{d\mid m}\operatorname{odd}(|\widetilde{V}_{d}|)$ divides $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|$, where $\widetilde{V}_{d}$ is as defined in Definition 10.16. It therefore suffices to prove that, for each $m$ listed above, we have $\operatorname{odd}(|\widetilde{V}_{d}|)\neq 1$ for some $d\mid m$.
@@
In each case, $\operatorname{odd}(c_{2p})\neq 1$ and so $\operatorname{odd}(|\widetilde{V}_{2p}|)\neq 1$. Hence $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|\neq 1$ for $m=2p$ where $p\in\{11,13,17,19,29\}$.
@@
In the cases $(p,q)\in\{(3,11),(5,7),(3,13)\}$, $\operatorname{odd}(c_{pq})\neq 1$ and so $\operatorname{odd}(|\widetilde{V}_{pq}|)\neq 1$. Hence $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|\neq 1$ for $m=pq$ or $2pq$ where $(p,q)\in\{(3,11),(5,7),(3,13)\}$.
We deal with the three remaining cases $m=21,30$, $42$ directly from the definition of $\widetilde{V}_{m}$:
@@
and $E=\langle\zeta_{15},\mathbb{Z}[\lambda_{15}]^{\times}\rangle$ has index two in $\mathbb{Z}[\zeta_{15}]^{\times}$. By the same argument as in Proposition 10.23, we get that $|\widetilde{V}_{30}|=|\operatorname{coker}(\Psi_{30}^{+})|$ is divisible by
@@
since $f_{2}=\operatorname{ord}_{15}(2)=4$ and $g_{2}=1$. Since $\operatorname{odd}(c_{30})\neq 1$, this implies that $\operatorname{odd}(|\widetilde{V}_{30}|)\neq 1$. Hence $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{15})\}|\neq 1$.
We deal with the remaining cases $m=21$, $42$ by computing the involution on $\widetilde{V}_{21}$ explicitly. This turns out to be necessary since, by Remark 11.11, we have $|D(\mathbb{Z}C_{m})|=4$ in each case and so $\operatorname{odd}(|\widetilde{V}_{21}|)=\operatorname{odd}(|\widetilde{V}_{42}|)=1$. If $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$ for $m=21$ or $42$ then, by Proposition 10.18, we would have $|2\cdot\widetilde{V}_{21}|=1$. Hence it suffices to prove that $|2\cdot\widetilde{V}_{21}|\neq 1$.
We now claim that $\widetilde{V}_{21}=\operatorname{coker}(\Psi_{21}^{+})\cong\mathbb{Z}/4$, which implies that $2\cdot\widetilde{V}_{21}\cong\mathbb{Z}/2$. Firstly, by the proof of Proposition 10.23, we have that
@@
where $1\in\mathbb{Z}/21$ maps to $\Psi_{21}^{+}(\zeta_{21})$ and $1\in\mathbb{Z}/2$ maps to $\Psi_{21}^{+}(1-\zeta_{21})$.
Since $g_{3}=1$, $\mathbb{F}_{3}[\zeta_{7}]\cong\mathbb{F}_{3}(\zeta_{7})$ is a field and $\mathbb{F}_{3}[\lambda_{7}]\cong\mathbb{F}_{3}(\lambda_{7})$ is a subfield. This implies that $\frac{\mathbb{F}_{3}[\zeta_{7}]^{\times}}{\mathbb{F}_{3}[\lambda_{7}]^{\times}}\cong\frac{\mathbb{Z}/(3^{6}-1)}{\mathbb{Z}/(3^{3}-1)}\cong\mathbb{Z}/28$. We have $\mathbb{F}_{7}[\zeta_{3}]=\mathbb{Z}[x]/\langle 7,1+x+x^{3}\rangle=\mathbb{Z}[x]/\langle 7,(x-2)(x-4)\rangle\cong\mathbb{F}_{7}\times\mathbb{F}_{7}$ where $\zeta_{3}\mapsto(2,4)$ and so $\mathbb{F}_{7}[\zeta_{3}]^{\times}\cong(\mathbb{Z}/6)^{2}$. Since $\mathbb{F}_{7}[\lambda_{3}]^{\times}=\mathbb{F}_{7}^{\times}\cong\mathbb{Z}/6$, this implies that $\frac{\mathbb{F}_{7}[\zeta_{3}]^{\times}}{\mathbb{F}_{7}[\lambda_{3}]^{\times}}\cong\mathbb{Z}/6$. Hence $D:=\frac{\mathbb{F}_{3}[\zeta_{7}]^{\times}}{\mathbb{F}_{3}[\lambda_{7}]^{\times}}\oplus\frac{\mathbb{F}_{7}[\zeta_{3}]^{\times}}{\mathbb{F}_{7}[\lambda_{3}]^{\times}}\cong\mathbb{Z}/28\oplus\mathbb{Z}/6$.
Now $\Psi_{21}^{+}(\zeta_{21})=[(\zeta_{7},\zeta_{3})]$ where $\zeta_{7}]\in\mathbb{Z}/28$ has order $7$ and $\zeta_{3}]\in\mathbb{Z}/6$ has order $3$. This implies that $\operatorname{coker}(\mathbb{Z}/21\to D)\cong\mathbb{Z}/4\oplus\mathbb{Z}/2$. Note that $\Psi_{21}^{+}(1-\zeta_{21})=[(1-\zeta_{7},1-\zeta_{3})]$. The isomorphism $\mathbb{F}_{7}[\zeta_{3}]^{\times}/\mathbb{F}_{7}[\lambda_{3}]^{\times}\to\mathbb{Z}/6$ sends $1-\zeta_{3}\mapsto 1-2=-1$ and so the image of $\Psi_{21}^{+}(1-\zeta_{21})$ in $\operatorname{coker}(\mathbb{Z}/21\to D)\cong\mathbb{Z}/4\oplus\mathbb{Z}/2$ has the form (,1)(*,-1). Since it has order two, this implies that $\operatorname{coker}(\Psi_{21}^{+})\cong\operatorname{coker}(\mathbb{Z}/21\oplus\mathbb{Z}/2\to D)\cong\mathbb{Z}/4$ as required. ∎
We now prove (A3). Our approach is to use Lemma 11.3 (iv), and is analogous to our proof of (A2). We begin by classifying the $m\geq 2$ square-free for which $\prod_{d\mid m}|\{x\in C(\mathbb{Z}[\zeta_{d}])\mid x=-\overline{x}\}|=1$. We then determine the subset of these values for which $|\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}|=1$.
@@
If $h_{m}^{-}=1$, then $h_{d}^{-}=1$ for all $d\mid m$ and so $\prod_{d\mid m}|\{x\in C(\mathbb{Z}[\zeta_{d}])\mid x=-\overline{x}\}|=1$. Conversely, if $h_{m}^{-}\neq 1$ and $m\not\in\{29,39,58,78\}$, then Proposition 11.5 implies that $|\{x-\overline{x}\mid x\in C(\mathbb{Z}[\zeta_{m}])\}|\neq 1$ and so $|\{x\in C(\mathbb{Z}[\zeta_{m}])\mid x=-\overline{x}\}|\neq 1$. It now suffices to show that, if $m\in\{29,39,58,78\}$, then $|\{x\in C(\mathbb{Z}[\zeta_{m}])\mid x=-\overline{x}\}|\neq 1$.
@@
We now prove the following. By Lemma 11.3 (iv), this completes the proof of (A3).
@@
If $|\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}|=1$, then $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$. By Propositions 10.6, 11.6, LABEL:, and 11.8, the $m\geq 2$ square-free for which $\prod_{d\mid m}|\{x\in C(\mathbb{Z}[\zeta_{d}])\mid x=-\overline{x}\}|=1$ and $|\{x-\overline{x}\mid x\in D(\mathbb{Z}C_{m})\}|=1$ are as follows:
@@
If $m=p$ for $p\leq 17$ prime or $m=2p$ for $p\in\{3,5,7\}$, then Lemma 11.7 implies that $D(\mathbb{Z}C_{m})=0$ and so $|\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}|=1$. If $m=15$, then it is shown in [CN72, p. 48] that $|D(\mathbb{Z}C_{15})|=2$. This implies that $D(\mathbb{Z}C_{15})\cong\mathbb{Z}/2$ has the trivial action and so $\{x\in D(\mathbb{Z}C_{m})\mid x=-\overline{x}\}\cong\mathbb{Z}/2$. ∎
Finally, we prove (A4). These results are implied by computations used in the proofs of Propositions 11.6 and 11.9, but we repeat them here for the convenience of the reader.
@@
This completes the proofs of (A1)-(A4) and so, by the discussion at the start of this section, completes the proof of Theorem 11.1.
@@
$A_{m}:=\frac{\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid x=-\overline{x}\}}{\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}}\cong\widehat{H}^{1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{m}))$
where the isomorphism comes from Proposition 9.4. Similarly to the proof of Theorem 11.1, we begin by noting that it now suffices to prove the following two statements.
@@
To see this, note that (B1) directly implies Theorem 11.2 (i) since $h_{3}=1$ is odd (see, for example, Proposition 10.6). Next, Theorem 11.2 (ii) follows from (B2) since it implies that
$\sup_{k\leq m}|A_{2^{k}}|$ $\geq\max\{|A_{2^{k}}|,|A_{2^{k-1}}|\}\geq\sqrt{|A_{2^{k}}|\cdot|A_{2^{k-1}}|}$
$\geq 2^{2^{k-3}-1}\geq 2^{2^{m-3}-1}$
@@
We again make use of the following short exact sequence of $\mathbb{Z}C_{2}$-modules (see Section 10.1):
@@
where $D(\mathbb{Z}C_{m})$ has the induced involution, and each $C(\mathbb{Z}[\zeta_{d}])$ has the involution induced by conjugation. Each of statements (B1) and (B2) are proven via the following lemma.
@@
$\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{m})_{2})$$A_{m}$$\bigoplus_{d\mid m}\widehat{H}^{1}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{2})$$\bigoplus_{d\mid m}\widehat{H}^{0}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{2})$$\widehat{H}^{0}(C_{2};C(\mathbb{Z}C_{m})_{2})$$\widehat{H}^{0}(C_{2};D(\mathbb{Z}C_{m})_{2}).$$\scriptstyle{\partial}\scriptstyle{\partial}$
@@
    If $h_{m}$ is odd, then $A_{m}\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{m})_{2})$
@@
    If $|D(\mathbb{Z}C_{m})|$ is odd, then $A_{m}\cong\bigoplus_{d\mid m}\widehat{H}^{1}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{2})$.
@@
By Proposition 9.7, we have that $A_{m}\cong\widehat{H}^{1}(C_{2};C(\mathbb{Z}C_{m})_{2})$. The short exact sequence stated above induces a short exact sequence on their $2$-primary submodules:
$0\to D(\mathbb{Z}C_{m})_{2}\to C(\mathbb{Z}C_{m})_{2}\to\bigoplus_{d\mid m}C(\mathbb{Z}[\zeta_{d}])_{2}\to 0.$
@@
$\widehat{H}^{n}(C_{2};\bigoplus_{d\mid m}C(\mathbb{Z}[\zeta_{d}])_{2})\cong\bigoplus_{d\mid m}\widehat{H}^{n}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{2})$
@@
To prove (i), suppose $h_{m}$ is odd. Then $h_{d}$ is odd for all $d\mid m$ since $h_{d}\mid h_{m}$ [Wa97, p. 205]. This implies that $C(\mathbb{Z}[\zeta_{d}])_{2}=0$ for all $d\mid m$ and so $\bigoplus_{d\mid m}\widehat{H}^{n}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{2})=0$ for all $n\in\mathbb{Z}$. The 6-periodic exact sequence then gives that $A_{m}\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{m})_{2})$.
To prove (ii), suppose $|D(\mathbb{Z}C_{m})|$ is odd. Then $D(\mathbb{Z}C_{m})_{2}=0$, $\widehat{H}^{n}(C_{2};D(\mathbb{Z}C_{m})_{2})=0$ for all $n$, so the 6-periodic exact sequence gives that $A_{m}\cong\bigoplus_{d\mid m}\widehat{H}^{1}(C_{2};C(\mathbb{Z}[\zeta_{d}])_{2})$. If $h_{m}$ and $|D(\mathbb{Z}C_{m})|$ are odd, then $D(\mathbb{Z}C_{m})_{2}=0$ and so $A_{m}\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{m})_{2})=0$, which proves (iii). ∎
We now prove (B1).
@@
We now prove (B2). By the discussion above, this completes the proof of Theorem 11.2.
@@
$A_{2^{n}}\cong\widehat{H}^{1}(C_{2};D(\mathbb{Z}C_{2^{n}})_{2}).$
By Propositions 9.5 and 10.20, we have a $6$-periodic exact sequence of finite abelian groups:
$\widehat{H}^{1}(C_{2};V_{2^{n+1}})$$A_{2^{n+1}}$$A_{2^{n}}$$\widehat{H}^{0}(C_{2};D(\mathbb{Z}C_{2^{n}}))$$\widehat{H}^{0}(C_{2};D(\mathbb{Z}C_{2^{n+1}}))$$\widehat{H}^{0}(C_{2};V_{2^{n+1}}).$$\scriptstyle{\partial}\scriptstyle{\partial}$